Prev Next

SW&IT Meeting Minutes 2018-10-22

Attendees: David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

Previous Action Items

  • xxx

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) PrimarySorted ascending Secondary Week
particle monitors AC Petr 4
AzCam
AC (2)
Petr (2)
2
Weatherstation
AC (2)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
OAC
AC (5)
Petr (4)
1
DIMM
AC (5)
Petr (2)
2
INDI Igor (2)
Matthieu (2)
3
ALH Igor(2)
Matthieu (2)
1
GCS
Petr (5)
Matthieu (2)
1
Telemetry
Petr (5)
Matthieu (2)
2
OVMS
Petr (5)
Matthieu (2)
3
Telemetry cleanup Stephen (1) Petr 3
Status Server Stephen (2)
Matthieu (2)
2
Telem-VIZ
Yang (1)
Stephen (1) 3
MODS
Yang (2)
Igor (1/2) 2
LBC
Yang (3)
Igor(2) 2

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • LUCI1 IP address
  • 4D clone - Xianju leading, discussed some steps, are we all clear what this is all about? XZ to summarize them (4D install on Win10/7, etc)
  • FOS Wiki - ingest ARGOS/Arcetri (DONE)
  • Magic laptop clone/vm - news from Microgate? (Got the change IP linux based client from Microgate - in rcs? how about the remaining eng. GUI interface?)
  • SOUL/FLAO Oct18 in lab then telescope
  • ARGOS/ci - add test, re-enable 'monit'
  • ARGOS deploy at telescope (Gustavo email)
  • Tucson CR, UPS/power plan (harware purchase before end of the year)
  • LBTI/shark+ilocater IIF/INDI (comments asked, XZ?)
  • Vsphere upgrade, capture in plan
  • Luci web-io cold heads state in telemetry for vibration
  • Luci H/O doc comments from RR
  • Tucson CR power upgrade - make plans now

This week in github

Team reports

PK

  • finished LBC NS tracking bug fix
  • some collimation look
  • PMCL troubleshooting & repairs
  • GCS WFS thread redesign
next:

AC

Last week
  • Vacation on Monday
  • HO tech talks: OAC and DIMM
    • Uploaded slides to wiki
  • Rename oac-tucson to oac-dev
    • SH strongly recommends NOT changing the name
  • ARGOS Docker
    • monit is failing

This week
  • Check AGW1 motions and cameras
  • Verify ARGOS deployments on obs machines with Gustavo
    • run simulator
  • ARGOS internal meeting on Wednesday
  • AROGS Docker
  • svn2git: oac, dimm?

SH

Last week:
  • ARGOS wiki import
  • LBTI host/VLAN changes for Paul G.
  • Replace Xianyu's machine
  • Virtualize abell.as.arizona.edu
  • Docker test environment
  • VM for Yang
This week:
  • Talk to John H., checkout abell, build FC kickstart
  • AGW1 onto telescope network changes
  • Utility building network changes
  • Dockerize "analysis" and move machine to Core-OS?
  • Bubble off pdb? Still looking for replacement pdb.
  • Finished copy of "AO" machines to USB. What next?

YZ

  • Web front programming learning and prototyping:
    • Setup a Fedora 28 VM with Stephen (ao-cravat)
    • Google Polymer 3 + Web components learning and prototyping.
  • Final meeting with Kellee on LBC and Misc things.
  • LBC support wiki page editing (with Igor).
  • Some work on collimation data (maybe more to come).
  • Old SVN tools/python code now on Github.

MB

  • management OC, WPC
  • jenkins-ci: ddclient, irc, facum, weather, ds9
  • svn2git: python-lbto, ddclient, facsum
  • h52csv reimplementation in python (x2 speed, x 0.1 code)
  • discuss PBP, 4D with Xianyu
  • discuss JH's abell transition
  • look at misc Wiki formatting issues
  • litHtml continued
  • sci-arch meeting
next:
  • svn2git: last push
  • skytech LBC update Fernando will give and update this week
  • OVMS/TT network in AO cabinet
  • SHARK-VIS Indi/IIF interface
  • new candidate phone interviews, post ad on select newsgroup
  • sci-arch follow up
  • Tucson power way forward with DE
  • Obs-ws way forward with DH

XZ

  • Fix SX and DX AdSecs starting issue
  • IDL license and license server installation (back and forth) on SX and DX AdSec with Stephen and Matthieu
  • SX and DX AdSec network renaming with Stephen
  • PBP dicussion and setup lab with final board (get some BCU readout)
  • 4D interferometer diucssion with Matthieu, AO, and 4D
  • changed the BCU IP with a small programm
  • Setup the 4D interferometer in the lab with IT
next:
  • Working on PBP lab setup
  • run different 4D interferometer SW with different OS systems
  • support AO operation

DE

Last week:
  • AGw work
    • Camera controller cards not working. Swapping them in the PCI slots fixed the issue, but it's not clear why
    • Jonathan is very willing to push all of these "computer" issues off on us; where is the line at which we stop?
  • Sketch of server room power, will digitize
  • Discovered major showstopper with Synology's cloud backup program
    • They rewrite local, production filenames to prevent case conflicts on the remote filesystem even if the remote is case-aware
    • I've stopped the Synology backups of the archive and restored all filenames to original--I do not think anyone was affected before I fixed it.
    • Moving to a slower but respectful method using rsync
  • Set up Windows 10 test VM for controller software installation tests
    • 150.135.245.159
    • RDP is open: U: lbto P: aotest
    • Shall I rip the install CD and make it available to the VM?
  • Synology Volume 3 rebuild. 2 disks at once took ~1 hour per percent completion, but it's finished now and back to full double-parity protection.
This week:
  • Vacation Friday
  • Send new DIN-mount switch to mountain
  • Add detail to power map while digitizing
  • Andrew: AGw fragility testing re: card slot dependency?

IS

Last Week
  • Documentation
    • INDI/ALH/LBC/MODS
  • OT
    • MODs script bug fix

This week
  • Wrapping up Kellee's handover
  • LUCI AFC testing/support
  • OT AFC development?
Topic revision: r14 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback